Three times the measure of a complement of a certain angle is equal to 30 more than the measure of a supplement of that angle. Find the measure of each angle.
Answer by Boreal(15235)

x=angle
90-x=complement
3(90-x)=30+(180-x), where 180-x is the supplement of the angle
270-3x=210-x
60=2x
x=30 deg, the angle
complement is 60 deg
3*60-30=supplement=150
supplement is 150 deg